OpenStudy (anonymous):

Determine whether each expression is a polynomial? will write in a second ... soo i know whats a polynomial its an expression more than one term except i want to clarified that fractions and radicals make it not a polynomial

OpenStudy (marissalovescats):

Yes you are correct. Polynomials have more than 2 terms. Monomial -1 Binomial-2 Polynomial- 3+

OpenStudy (anonymous):

1) \[3x+\frac{ 4 }{ 3 }x^3-7\] 2) \[-5x^3+x-\frac{ 3 }{ x }\] 3) \[x^3\sqrt{5}+x-3\] 4) \[25x^2-x^3+x^5\]

OpenStudy (marissalovescats):

Which ones do you think are polynomials?

OpenStudy (anonymous):

i know they all have more than one term BUT would the radical make it not a polynomial

OpenStudy (marissalovescats):

Yes it would make it not

OpenStudy (anonymous):

the only one which i believe is a polynomial is the last one

OpenStudy (marissalovescats):

The last one yes. But there is one more

OpenStudy (anonymous):

would it be the first one

OpenStudy (anonymous):

because the second one has 3/x

OpenStudy (marissalovescats):

Fractions: